OpenStudy (anonymous):

Factor by grouping : x^3-4x^2+3x-12

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

x= x^2(x - 4) + 3(x - 4) = (x^2 + 3)(x - 4)

OpenStudy (skullpatrol):

Correct :)

OpenStudy (anonymous):

Yeah you're right

OpenStudy (skullpatrol):

$$ x^3-4x^2+3x-12 =x^2(x - 4) + 3(x - 4) = (x^2 + 3)(x - 4)$$